OpenStudy (anonymous):

Three point-charges of +2x10^-6 C, -3x10^-6 C and -3x10^-6 C are kept at the vertices A,B and C respectively of an equilateral triangle of side 20 cm. What should be the sign and magnitude of a charge to be placed at the mid-point M of side BC so that the charge at A remains in equillibrium??

OpenStudy (mani_jha):

 See the diagram. Let its charge be Q. It is acted upon by three forces, one from A, one from B, and one from C. Calculate force from B, C, using Coulomb's formula. r=20/2=10=0.1m(Taking Q as unknown charge) Now, add these forces. Now, calculate force from A(r=20/ So, Force from B must cancel the force from C. So, consider only the force due to A.